# Model-Agnostic Bounds for Augmented Inverse Probability Weighted Estimators’ Wald-Confidence Interval Coverage in Randomized Controlled Trials

### v1

@@ 6 Discussion
$\tilde{\mathcal{R}}(\delta,\delta^{\prime})$ $:=\frac{{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}}{4\delta^{2}M^{2}}+{\mathcal{C}}\left\{\frac{16M^{2}(1-\tau_{\pi})^{2}z_{\alpha}^{2}}{\tau_{\pi}^{2}\sigma_{\#,a}^{2}\sigma_{\dagger,a}^{2}}{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}\right\}^{1/3}$
$\quad+{\mathcal{C}}\frac{(F^{{\mathcal{T}}_{a}^{2}})^{2}J^{2}(2,{\mathcal{F}},M)({\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2})^{2}}{\sigma_{\#,a}^{4}(\delta^{\prime})^{4}M^{4}n}.$
Additionally under subgaussian conditions 10 and S2, if $2M\min\{\delta,\delta^{\prime}\}>{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}$, then (S3) holds with $\tilde{\mathcal{R}}(\delta,\delta^{\prime})$ replaced by
$\tilde{\mathcal{R}}^{\prime}(\delta,\delta^{\prime})$ $:=\exp\left\{-\frac{(2\delta M-{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2})^{2}}{2\tilde{s}_{1}^{2}{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}}\right\}+{\mathcal{C}}\frac{\tilde{s}_{3}M(1-\tau_{\pi})|z_{\alpha}|}{\tau_{\pi}\sigma_{\#,a}\sigma_{\dagger,a}}\sqrt{q{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}\log n}$
$\qquad+{\mathcal{C}}\frac{(F^{{\mathcal{T}}_{a}^{2}})^{2}J^{2}(2,{\mathcal{F}},M)}{\sigma_{\#,a}^{4}n}\exp\left\{-\frac{(2\delta^{\prime}M-{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2})^{2}}{\tilde{s}_{1}^{2}{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}}\right\}.$
Under Condition 11, the term $\{(1-\tau_{\pi})/\tau_{\pi}\}^{2}\|Q_{\#,a}-Q_{*,a}\|_{2}^{2}({\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2})^{2}$ can de dropped from the right-hand side of (S3).
The next theorem is a non-asymptotic version of Theorem 4 and Part 2 of Proposition 1. It implies Theorem 4 by taking $\delta=1$ and noting that ${\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}\leq\sqrt{{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}}$. The ${\mathrm{O}}(n^{-1/2})$ term might be able to be improved to ${\mathrm{o}}(n^{-1/2})$ or a more precise rate depending on the covering number of $\mathcal{F}$}, by taking $\delta\to 0$ at a suitable rate. Part 2 of Proposition 1 follows by noting that $\mathscr{B}$ in Theorem S4 is ${\mathrm{O}}(n^{-1/2})$ and Lemma S13.
@@
Under Conditions 1, 3, 4 and 9, for any $\delta>0$,
@@
$-\left[P_{*}\left\{\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}{\mathbb{E}}_{\tilde{Q}_{a}}[(\tilde{Q}_{a}-Q_{\#,a})^{2}]\right\}+2P_{*}\left\{\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}(Q_{\#,a}-Q_{*,a}){\mathbb{E}}_{\tilde{Q}_{a}}[\tilde{Q}_{a}-Q_{\#,a}]\right\}-{\mathrm{Var}}(\tilde{\psi}_{a})\right]\Bigg|$
@@
$\mathscr{B}:={\mathcal{C}}\frac{F^{{\mathcal{T}}_{a}^{2}}J(2\delta,{\mathcal{F}},M)}{\sqrt{n}}+{\mathcal{C}}\frac{F^{{\mathcal{T}}_{a}^{2}}J^{2}(2\delta,{\mathcal{F}},M)}{\delta^{2}n}+{\mathcal{C}}\frac{F^{{\mathcal{T}}_{a}^{2}}J(2,{\mathcal{F}},M){\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}}{\sigma_{\#,a}\delta^{2}M^{2}\sqrt{n}}.$
@@
$|\sigma_{\dagger,a}-\sigma_{\#,a}|$ $\leq\frac{\mathscr{B}}{\sigma_{\#,a}}+\frac{1-\tau_{\pi}}{\tau_{\pi}\sigma_{\#,a}}{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}+2\frac{1-\tau_{\pi}}{\tau_{\pi}\sigma_{\#,a}}\|Q_{\#,a}-Q_{*,a}\|_{2}{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}$
$\quad+\frac{{\mathcal{C}}}{\sigma_{\#,a}n}\left\{\sigma_{\#,a}^{2}+(F^{H_{a}})^{2}J^{2}(2,{\mathcal{F}},M)\right\}.$ (S5)
Under Condition 11, the terms $2P_{*}\left\{\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}(Q_{\#,a}-Q_{*,a}){\mathbb{E}}_{\tilde{Q}_{a}}[\tilde{Q}_{a}-Q_{\#,a}]\right\}$ and $2\frac{1-\tau_{\pi}}{\tau_{\pi}\sigma_{\#,a}}\|Q_{\#,a}-Q_{*,a}\|_{2}{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}$ can be dropped from (S4) and (S5), respectively. Under subgaussian condition 10, if $2\delta M>{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}$, the third term ${\mathcal{C}}\frac{F^{{\mathcal{T}}_{a}^{2}}J(2,{\mathcal{F}},M){\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}}{\delta^{2}M^{2}\sqrt{n}}$ in $\mathscr{B}$ can be replaced by
@@
$\frac{{\mathcal{C}}F^{{\mathcal{T}}_{a}^{2}}J(2,{\mathcal{F}},M)}{\sigma_{\#,a}\sqrt{n}}\exp\left\{-\frac{(2\delta M-{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2})^{2}}{2\tilde{s}_{1}^{2}{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}}\right\}.$
S3 Additional simulation details
@@
S4 Proof
S4.1 General preliminary lemmas
@@
$\Pr(\sqrt{n}P_{n}D_{a}(Q_{\#,a},\psi_{*,a})\leq z_{\alpha}\sigma_{\dagger,a}-\eta)-\Pr(|\sqrt{n}\{\tilde{\psi}_{a}-\psi_{*,a}-P_{n}D_{a}(Q_{\#,a},\psi_{*,a})\}-z_{\alpha}(\tilde{\sigma}_{a}-\sigma_{\dagger,a})|>\eta)$
$\leq\Pr(\sqrt{n}(\tilde{\psi}_{a}-\psi_{*,a})\leq z_{\alpha}\tilde{\sigma}_{a})$
$\leq\Pr(\sqrt{n}P_{n}D_{a}(Q_{\#,a},\psi_{*,a})\leq z_{\alpha}\sigma_{\dagger,a}+\eta)+\Pr(|\sqrt{n}\{\tilde{\psi}_{a}-\psi_{*,a}-P_{n}D_{a}(Q_{\#,a},\psi_{*,a})\}-z_{\alpha}(\tilde{\sigma}_{a}-\sigma_{\dagger,a})|>\eta)$
@@
$\Pr(\sqrt{n}P_{n}D_{a}(Q_{\#,a},\psi_{*,a})\leq z_{\alpha}\sigma_{\dagger,a}+\eta)$ $\leq 1-\alpha+{\mathcal{C}}\frac{\rho_{\#,a}}{\sigma_{\#,a}^{3}\sqrt{n}}+\phi(z_{\alpha})z_{\alpha}\frac{\sigma_{\dagger,a}-\sigma_{\#,a}}{\sigma_{\#,a}}+{\mathcal{C}}\frac{(\sigma_{\dagger,a}-\sigma_{\#,a})^{2}}{\sigma_{\#,a}^{2}}+{\mathcal{C}}\frac{\eta}{\sigma_{\#,a}},$
$\Pr(\sqrt{n}P_{n}D_{a}(Q_{\#,a},\psi_{*,a})\leq z_{\alpha}\sigma_{\dagger,a}-\eta)$ $\geq 1-\alpha-{\mathcal{C}}\frac{\rho_{\#,a}}{\sigma_{\#,a}^{3}\sqrt{n}}+\phi(z_{\alpha})z_{\alpha}\frac{\sigma_{\dagger,a}-\sigma_{\#,a}}{\sigma_{\#,a}}-{\mathcal{C}}\frac{(\sigma_{\dagger,a}-\sigma_{\#,a})^{2}}{\sigma_{\#,a}^{2}}-{\mathcal{C}}\frac{\eta}{\sigma_{\#,a}}.$
@@
$\Pr(\sqrt{n}P_{n}D_{a}(Q_{\#,a},\psi_{*,a})\leq z_{\alpha}\sigma_{\dagger,a}+\eta)$
$=\Pr\left(\sqrt{n}\frac{P_{n}D_{a}(Q_{\#,a},\psi_{*,a})}{\sigma_{\#,a}}\leq z_{\alpha}+\frac{z_{\alpha}(\sigma_{\dagger,a}-\sigma_{\#,a})+\eta}{\sigma_{\#,a}}\right)$
@@
The next lemma provides the first and second moments of $H_{a}(Q)$ (defined in Section 2).
@@
$P_{*}H_{a}(Q)$ $={\mathbb{E}}\left[\left(1-\frac{{\mathbbm{1}}(A=a)}{\pi_{*}(a\mid X)}\right)\left\{Q(X)-Q_{\#,a}(X)\right\}\right]$
$={\mathbb{E}}\left[{\mathbb{E}}\left[\left(1-\frac{{\mathbbm{1}}(A=a)}{\pi_{*}(a\mid X)}\right)\left\{Q(X)-Q_{\#,a}(X)\right\}\mid X\right]\right]=0,$
$P_{*}H_{a}(Q)^{2}$ $={\mathbb{E}}\left[\left(1-\frac{{\mathbbm{1}}(A=a)}{\pi_{*}(a\mid X)}\right)^{2}\left\{Q(X)-Q_{\#,a}(X)\right\}^{2}\right]$
$={\mathbb{E}}\left[{\mathbb{E}}\left[\left(1-\frac{{\mathbbm{1}}(A=a)}{\pi_{*}(a\mid X)}\right)^{2}\left\{Q(X)-Q_{\#,a}(X)\right\}^{2}\mid X\right]\right]$
$={\mathbb{E}}\left[{\mathbb{E}}\left[\left(1-2\frac{{\mathbbm{1}}(A=a)}{\pi_{*}(a\mid X)}+\frac{{\mathbbm{1}}(A=a)}{\pi_{*}(a\mid X)^{2}}\right)\left\{Q(X)-Q_{\#,a}(X)\right\}^{2}\mid X\right]\right]$
$={\mathbb{E}}\left[\left(1-2+\frac{1}{\pi_{*}(a\mid X)}\right)\left\{Q(X)-Q_{\#,a}(X)\right\}^{2}\right]$
$={\mathbb{E}}\left[\frac{1-\pi_{*}(a\mid X)}{\pi_{*}(a\mid X)}\left\{Q(X)-Q_{\#,a}(X)\right\}^{2}\right].$
@@
$\quad+\left\{\frac{{\mathbbm{1}}(A=a)}{\pi_{*}(a\mid X)^{2}}-2\frac{{\mathbbm{1}}(A=a)}{\pi_{*}(a\mid X)}+1\right\}Q(X)^{2}.$
@@
${\mathbb{E}}_{\hat{Q}}[{\mathcal{T}}_{a}(\hat{Q})(V)^{2}]$ $=\frac{{\mathbbm{1}}(A=a)}{\pi_{*}(a\mid X)^{2}}Y^{2}-2\frac{{\mathbbm{1}}(A=a)(1-\pi_{*}(a\mid X))}{\pi_{*}(a\mid X)^{2}}Y{\mathbb{E}}_{\hat{Q}}[\hat{Q}(X)]$
$\quad+\left\{\frac{{\mathbbm{1}}(A=a)}{\pi_{*}(a\mid X)^{2}}-2\frac{{\mathbbm{1}}(A=a)}{\pi_{*}(a\mid X)}+1\right\}{\mathbb{E}}_{\hat{Q}}[\hat{Q}(X)^{2}],$
${\mathbb{E}}\left[P_{*}\{{\mathcal{T}}_{a}(\hat{Q})^{2}-{\mathcal{T}}_{a}(Q_{\#,a})^{2}\}\right]$ $={\mathbb{E}}\left[P_{*}\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}(\hat{Q}-Q_{\#,a})^{2}\right]$
$\quad+2{\mathbb{E}}\left[P_{*}\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}(\hat{Q}-Q_{\#,a})(Q_{\#,a}-Q_{*,a})\right],$
$P_{*}\{{\mathcal{T}}_{a}(\hat{Q})^{2}-{\mathbb{E}}_{\hat{Q}}[{\mathcal{T}}_{a}(\hat{Q})^{2}]\}$ $=P_{*}\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}\left\{-2Q_{*,a}(\hat{Q}-{\mathbb{E}}_{\hat{Q}}[\hat{Q}])+\hat{Q}^{2}-{\mathbb{E}}_{\hat{Q}}[\hat{Q}^{2}]\right\}.$

### v2

@@ 6 Discussion
[…]
$\quad+{\mathcal{C}}\frac{(F^{{\mathcal{T}}_{a}^{2}})^{2}J^{2}(2,{\mathcal{F}},M){\color[rgb]{0,0,1}({E}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2})^{2}}}{\sigma_{\#,a}^{4}(\delta^{\prime})^{4}M^{4}n}.$
Additionally under approximate sub-Weibull conditions 11 and 12, (S3) holds with $\tilde{\mathcal{R}}(\delta,\delta^{\prime})$ replaced by
$\tilde{\mathcal{R}}^{\prime}(\delta,\delta^{\prime})$ $:=\tilde{a}_{1}n^{\tilde{d}_{1}}\exp\left\{-\tilde{c}_{1}\left(\frac{2{\color[rgb]{0,0,1}\delta}M}{{\color[rgb]{0,0,1}\tilde{r}_{1}(n)}}\right)^{\tilde{q}_{1}}\right\}+\tilde{b}_{1}{\color[rgb]{0,0,1}n^{-1/2}\log n}+{\mathcal{C}}\frac{|z_{\alpha}|}{\sigma_{\#,a}\sigma_{\dagger,a}}{\color[rgb]{0,0,1}\tilde{r}_{2}(n)}\left(\frac{q+\tilde{d}_{2}}{\tilde{c}_{2}}{\color[rgb]{0,0,1}\log n}\right)^{1/\tilde{q}_{2}}$
$\qquad+{\mathcal{C}}\frac{(F^{{\mathcal{T}}_{a}^{2}})^{2}J^{2}(2,{\mathcal{F}},M)}{\sigma_{\#,a}^{4}{\color[rgb]{0,0,1}n}}\left[\tilde{a}_{1}^{2}\exp\left\{-2\tilde{c}_{1}\left(\frac{2\delta^{\prime}M}{{\color[rgb]{0,0,1}\tilde{r}_{1}(n)}}\right)^{\tilde{q}_{1}}\right\}+\tilde{b}_{1}^{2}{\color[rgb]{0,0,1}n^{-1}(\log n)^{2}}\right].$
Under Condition 13, the term $\{(1-\tau_{\pi})/\tau_{\pi}\}^{2}\|Q_{\#,a}-Q_{*,a}\|_{2}^{2}{\color[rgb]{0,0,1}({E}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2})^{2}}$ can de dropped from the right-hand side of (S3).
The next theorem is a non-asymptotic version of Theorem 4 and Part 2 of Proposition 1. It implies Theorem 4 by taking $\delta=1$ and noting that ${E}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}\leq\sqrt{{E}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}}$ and ${\mathrm{var}}(\tilde{\psi}_{a})={\mathrm{O}}(n^{-1})$ by Lemma S13 under the setting of Theorem 4. The ${\mathrm{O}}(n^{-1/2})$ term might be able to be improved to ${\mathrm{o}}(n^{-1/2})$ or a more precise rate depending on the covering number of $\mathcal{F}$}, by taking $\delta\to 0$ at a suitable rate. Part 2 of Proposition 1 follows by noting that $\mathscr{B}$ in Theorem S4 is ${\mathrm{O}}(n^{-1/2})$ and Lemma S13.
@@
Under Conditions 1, 3, 4 and 10, for any $\delta>0$,
@@
$-\left[P_{*}\left\{\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}{E}_{\tilde{Q}_{a}}[{\color[rgb]{0,0,1}(\tilde{Q}_{a}-Q_{\#,a})^{2}}]\right\}+2P_{*}\left\{\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}(Q_{\#,a}-Q_{*,a}){E}_{\tilde{Q}_{a}}[{\color[rgb]{0,0,1}\tilde{Q}_{a}-Q_{\#,a}}]\right\}-{\color[rgb]{0,0,1}{\mathrm{var}}(\tilde{\psi}_{a})}\right]\Bigg|$
@@
[…]
@@
[…]
$\quad+\frac{{\mathcal{C}}}{\sigma_{\#,a}{\color[rgb]{0,0,1}n}}\left\{\sigma_{\#,a}^{2}+(F^{H_{a}})^{2}J^{2}(2,{\mathcal{F}},M)\right\}.$ (S5)
[…]
@@
$\frac{{\mathcal{C}}F^{{\mathcal{T}}_{a}^{2}}J(2,{\mathcal{F}},M)}{\sigma_{\#,a}{\color[rgb]{0,0,1}\sqrt{n}}}\left[\tilde{a}_{1}{\color[rgb]{0,0,1}n^{\tilde{d}_{1}}}\exp\left\{-\tilde{c}_{1}\left(\frac{2{\color[rgb]{0,0,1}\delta}M}{{\color[rgb]{0,0,1}\tilde{r}_{1}(n)}}\right)^{\tilde{q}_{1}}\right\}+\tilde{b}_{1}{\color[rgb]{0,0,1}n^{-1/2}\log n}\right].$
S2 Series regression: an example of outcome model estimators with approximate sub-Weibull tails
In this section, I provide an example of outcome model estimators satisfying approximate sub-Weibull conditions with constants uniform over $n$. The bounds used may be overly conservative, so approximate sub-Weibull conditions may hold more generally. The notation introduced in this section is used only within this section. Consider observing an i.i.d. sample $(X_{i},Y_{i})_{i=1}^{n}$ and estimating $Q_{*}:x\mapsto{E}[Y\mid X=x]$ with series regression. Let $\{\phi_{k}\}_{k=1}^{\infty}$ be a basis of functions chosen by the user and $\Phi_{K}:x\mapsto(\phi_{1}(x),\ldots,\phi_{K}(x))^{\top}$ be the vector of the first $K$ basis functions. Usually, $K=K(n)$ grows to infinity at an appropriate rate slower than $n$. In the special case of ordinary linear regression, $K$ is a constant. The corresponding series estimator $\hat{Q}$ of $Q_{*}$ is $x\mapsto\Phi_{K}(x)^{\top}\hat{\beta}$, where $\hat{\beta}:=\{P_{n}\Phi_{K}(X)\Phi_{K}(X)^{\top}\}^{+}\{P_{n}\Phi_{K}(X)Y\}$ and the superscript “++” denotes the Moore-Penrose pseudo-inverse. Let $Q_{\#}:x\mapsto\Phi_{K}(x)^{\top}\beta_{\#}$ where $\beta_{\#}:=\{P_{*}\Phi_{K}(X)\Phi_{K}(X)^{\top}\}^{+}\{P_{*}\Phi_{K}(X)Y\}$, so $Q_{\#}$ is the true best approximation to $Q_{*}$ with the first $K$ basis functions. This scenario corresponds to estimating $Q_{*,a}$ with series regression among individuals with $A=a$.
Let $\lambda_{\max}$ and $\lambda_{\min}$ denote the maximum and minimum eigenvalues of a matrix, respectively. Let $\|\cdot\|_{{\ell_{2}}}$ denote the $\ell_{2}$ norm of a vector.
SeriesCond S1.
There exists a sequence of constants $\{\xi_{K}\}_{K=1}^{\infty}$ and constant invertible $K\times K$ matrices $\{T_{K}\}_{K=1}^{\infty}$ such that $\sup_{x\in\mathcal{X}}\|T_{K}\Phi_{K}(x)\|_{\ell_{2}}\leq\xi_{K}$.
SeriesCond S2.
There exist constants $0<\underline{\lambda}\leq\bar{\lambda}<\infty$ such that $\underline{\lambda}\leq\lambda_{\min}(P_{*}T_{K}\Phi_{K}(X)\Phi_{K}(X)^{\top}T_{K}^{\top})\leq\lambda_{\max}(P_{*}T_{K}\Phi_{K}(X)\Phi_{K}(X)^{\top}T_{K}^{\top})\leq\bar{\lambda}$ for all $K$.
SeriesCond S3.
The following rates hold: $K/n={\mathrm{o}}(1),\xi_{K}^{2}/K={\mathrm{O}}(1)$.
SeriesCond S4.
$Y$ is bounded and ${\mathrm{var}}(Y)\mid X$ is bounded away from zero.
These conditions are common in the series regression literature (e.g., Newey,, 1997). The first part of the growth rate condition S3 is very weak because only the variance $\hat{Q}-Q_{\#}$ needs to be controlled but the bias $Q_{\#}-Q_{*}$ need not. The second part can be restrictive, but it is satisfied by some common series, including trigonometric polynomials, splines, orthogonal wavelets (Chen,, 2007). The boundedness of $Y$ in SeriesCond S4 ensures sub-Gaussianity. It is generally mild and might be able to be relaxed to light-tail distributions.
Proposition S1.
There exist constants $a,b,c>0$ independent of $(K,n)$ such that, for all $t>0$,
${\mathrm{pr}}\left(\frac{\|\hat{Q}-Q_{\#}\|_{2}}{K/n}>t\right)\leq a\exp\left(-ct\right)+b\sqrt{\log n/n}.$
Here, $r(n)$ is the convergence rate of $\|\hat{Q}-Q_{\#}\|_{2}$. Since the sample size for estimating the outcome model has a tiny probability (decaying exponentially with $n$) to be insufficient, Conditions 7 and Condition 11 hold. By Lemma S6 and the boundedness of $\hat{Q}$, the tail of $P_{*}{\mathcal{T}}_{a}(\hat{Q})^{2}$ is similar to $\|\hat{Q}-Q_{\#,a}\|_{2}$. Thus, Conditions 8 and 12 also hold.
S3 Impact of the choice of $K$ on the rates in Theorem 1
For the cross-fit procedure, the choice of the number of folds $K$ in cross-fitting under exchangeable sample-splitting (Condition 6) is an interesting question. If the impact of $K$ on ${E}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}$ is ignored, Theorem 1 suggests choosing a small $K$ because the bound increases with $K$. However, a small $K$ would generally lead to a larger ${E}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}$ because of fewer training data. Without additional knowledge or assumptions on how ${E}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}$ varies with training sample size n(K1)/Kn(K-1)/K, it appears impossible to obtain concrete theoretical guidance on the choice of $K$. Suppose that ${E}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}={\mathrm{O}}(\{n(K-1)/K\}^{-r})$ for some constant $r>0$. For example, if $\hat{Q}_{k,a}$ is estimated in a parametric model, $r$ can be anticipated to be one; if the usual ${\mathrm{o}}_{p}(n^{-1/4})$-rate requirement on $\hat{Q}_{k,a}$ is achieved, $r$ can be anticipated to be larger than 1/2. The resulting rate in Theorem 1 is $n^{-r/3}K^{(r+2)/3}(K-1)^{-r/3}$, or $n^{-r/2}K^{(r+1)/2}(K-1)^{-r/2}$ (up to log factors) under approximate sub-Weibull conditions 7 and 8. Figure S1 illustrates the multiplicative factor involving $K$ in the rate for a range of $r$ and $K$, and suggests that a small $K$ is generally preferable.
This result differs from the practical recommendation in Phillips et al., (2023), where a larger $K$ like 20 is recommended in smaller samples. Because I use a crude union bound to analyze cross-fitting, the rate in Theorem 1 might be suboptimal regarding $K$. In particular, I might exaggerate the loss in the rate due to a large $K$, so the above analysis might not accurately characterize the impact of $K$ in cross-fitting. A sound theoretical justification on the choice of $K$ is still an open question.
Figure S1: The factor concerning $K$ in the cross-fit Wald-CI coverage’s convergence rate to the nominal coverage. The number of folds $K$ ranges over even numbers between 2 and 20. The rate $r$ ranges from 0.4 (slow) to 1 (parametric). Subfigures (A) and (B) are for the cases with and without approximate sub-Weibull conditions 7 and 8, respectively. The points are dodged horizontally.
S4 Additional simulation details
@@
S5 Proof
S5.1 General preliminary lemmas
@@
${\mathrm{pr}}(\sqrt{n}P_{n}D_{a}(Q_{\#,a},\psi_{*,a})\leq z_{\alpha}\sigma_{\dagger,a}-\eta)-{\mathrm{pr}}(|\sqrt{n}\{\tilde{\psi}_{a}-\psi_{*,a}-P_{n}D_{a}(Q_{\#,a},\psi_{*,a})\}-z_{\alpha}(\tilde{\sigma}_{a}-\sigma_{\dagger,a})|>\eta)$
$\leq{\mathrm{pr}}(\sqrt{n}(\tilde{\psi}_{a}-\psi_{*,a})\leq z_{\alpha}\tilde{\sigma}_{a})$
$\leq{\mathrm{pr}}(\sqrt{n}P_{n}D_{a}(Q_{\#,a},\psi_{*,a})\leq z_{\alpha}\sigma_{\dagger,a}+\eta)+{\mathrm{pr}}(|\sqrt{n}\{\tilde{\psi}_{a}-\psi_{*,a}-P_{n}D_{a}(Q_{\#,a},\psi_{*,a})\}-z_{\alpha}(\tilde{\sigma}_{a}-\sigma_{\dagger,a})|>\eta)$
@@
${\mathrm{pr}}(\sqrt{n}P_{n}D_{a}(Q_{\#,a},\psi_{*,a})\leq z_{\alpha}\sigma_{\dagger,a}+\eta)$ $\leq 1-\alpha+{\mathcal{C}}\frac{\rho_{\#,a}}{\sigma_{\#,a}^{3}\sqrt{n}}+\phi(z_{\alpha})z_{\alpha}\frac{\sigma_{\dagger,a}-\sigma_{\#,a}}{\sigma_{\#,a}}+{\mathcal{C}}\frac{(\sigma_{\dagger,a}-\sigma_{\#,a})^{2}}{\sigma_{\#,a}^{2}}+{\mathcal{C}}\frac{\eta}{\sigma_{\#,a}},$
${\mathrm{pr}}(\sqrt{n}P_{n}D_{a}(Q_{\#,a},\psi_{*,a})\leq z_{\alpha}\sigma_{\dagger,a}-\eta)$ $\geq 1-\alpha-{\mathcal{C}}\frac{\rho_{\#,a}}{\sigma_{\#,a}^{3}\sqrt{n}}+\phi(z_{\alpha})z_{\alpha}\frac{\sigma_{\dagger,a}-\sigma_{\#,a}}{\sigma_{\#,a}}-{\mathcal{C}}\frac{(\sigma_{\dagger,a}-\sigma_{\#,a})^{2}}{\sigma_{\#,a}^{2}}-{\mathcal{C}}\frac{\eta}{\sigma_{\#,a}}.$
@@
${\mathrm{pr}}(\sqrt{n}P_{n}D_{a}(Q_{\#,a},\psi_{*,a})\leq z_{\alpha}\sigma_{\dagger,a}+\eta)$
$={\mathrm{pr}}\left(\sqrt{n}\frac{P_{n}D_{a}(Q_{\#,a},\psi_{*,a})}{\sigma_{\#,a}}\leq z_{\alpha}+\frac{z_{\alpha}(\sigma_{\dagger,a}-\sigma_{\#,a})+\eta}{\sigma_{\#,a}}\right)$
@@
For any function $Q$ and $a^{\prime}\in\{0,1\}$, define function $H_{a^{\prime}}(Q):v=(x,a,y)\mapsto{\mathcal{T}}_{a^{\prime}}(Q)(v)-{\mathcal{T}}_{a^{\prime}}(Q_{\#,a^{\prime}})(v)=\{1-{\mathbbm{1}}(a=a^{\prime})/\pi_{*}(a^{\prime}\mid x)\}\{Q(x)-Q_{\#,a^{\prime}}(x)\}$. The next lemma provides the first and second moments of $H_{a}(Q)$.
@@
$P_{*}H_{a}(Q)$ $={E}\left[\left(1-\frac{{\mathbbm{1}}(A=a)}{\pi_{*}(a\mid X)}\right)\left\{Q(X)-Q_{\#,a}(X)\right\}\right]$
$={E}\left[{E}\left[\left(1-\frac{{\mathbbm{1}}(A=a)}{\pi_{*}(a\mid X)}\right)\left\{Q(X)-Q_{\#,a}(X)\right\}\mid X\right]\right]=0,$
$P_{*}H_{a}(Q)^{2}$ $={E}\left[\left(1-\frac{{\mathbbm{1}}(A=a)}{\pi_{*}(a\mid X)}\right)^{2}\left\{Q(X)-Q_{\#,a}(X)\right\}^{2}\right]$
$={E}\left[{E}\left[\left(1-\frac{{\mathbbm{1}}(A=a)}{\pi_{*}(a\mid X)}\right)^{2}\left\{Q(X)-Q_{\#,a}(X)\right\}^{2}\mid X\right]\right]$
$={E}\left[{E}\left[\left(1-2\frac{{\mathbbm{1}}(A=a)}{\pi_{*}(a\mid X)}+\frac{{\mathbbm{1}}(A=a)}{\pi_{*}(a\mid X)^{2}}\right)\left\{Q(X)-Q_{\#,a}(X)\right\}^{2}\mid X\right]\right]$
$={E}\left[\left(1-2+\frac{1}{\pi_{*}(a\mid X)}\right)\left\{Q(X)-Q_{\#,a}(X)\right\}^{2}\right]$
$={E}\left[\frac{1-\pi_{*}(a\mid X)}{\pi_{*}(a\mid X)}\left\{Q(X)-Q_{\#,a}(X)\right\}^{2}\right].$
@@
$\quad+\left\{\frac{{\mathbbm{1}}(A=a)}{\pi_{*}(a\mid X)^{2}}-2\frac{{\mathbbm{1}}(A=a)}{\pi_{*}(a\mid X)}+1\right\}Q(X)^{2},$
$P_{*}{\mathcal{T}}_{a}(Q)^{2}$ $=P_{*}\frac{{E}[Y^{2}\mid A=a,X=\cdot]}{\pi_{*}(a\mid\cdot)}-2P_{*}\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}Q_{*}Q+P_{*}\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}Q^{2}.$
@@
${E}_{\hat{Q}}[{\mathcal{T}}_{a}(\hat{Q})(V)^{2}]$ $=\frac{{\mathbbm{1}}(A=a)}{\pi_{*}(a\mid X)^{2}}Y^{2}-2\frac{{\mathbbm{1}}(A=a)(1-\pi_{*}(a\mid X))}{\pi_{*}(a\mid X)^{2}}Y{E}_{\hat{Q}}[\hat{Q}(X)]$
$\quad+\left\{\frac{{\mathbbm{1}}(A=a)}{\pi_{*}(a\mid X)^{2}}-2\frac{{\mathbbm{1}}(A=a)}{\pi_{*}(a\mid X)}+1\right\}{E}_{\hat{Q}}[\hat{Q}(X)^{2}],$
${E}\left[P_{*}\{{\mathcal{T}}_{a}(\hat{Q})^{2}-{\mathcal{T}}_{a}(Q_{\#,a})^{2}\}\right]$ $={E}\left[P_{*}\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}(\hat{Q}-Q_{\#,a})^{2}\right]$
$\quad+2{E}\left[P_{*}\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}(\hat{Q}-Q_{\#,a})(Q_{\#,a}-Q_{*,a})\right],$
$P_{*}\{{\mathcal{T}}_{a}(\hat{Q})^{2}-{E}_{\hat{Q}}[{\mathcal{T}}_{a}(\hat{Q})^{2}]\}$ $=P_{*}\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}\left\{-2Q_{*,a}(\hat{Q}-{E}_{\hat{Q}}[\hat{Q}])+\hat{Q}^{2}-{E}_{\hat{Q}}[\hat{Q}^{2}]\right\}.$